The next dividend payment by Hoffman, Inc., will be $2.50 per share. The dividends are anticipated to maintain a growth rate of 5.75 percent forever. Assume the stock currently sells for $48.60 per share.

What is the expected capital gains yield?
Answer #1

HI

Here next dividend D1 = $2.50

growth rate g= 5.75%

required return k = D1/P0 +g

= 2.5/48.60 + 0.0575

= 10.89%

So capital gin yield = (P1-P0)/P0

P1 = D2/(k-g)

P1 = D1*(1+g)/(k-g)

=2.5*(1+5.75%)/(10.89%-5.75%)

= 51.3945

Capital gain yield = (51.3945-48.60)/48.60

= 5.75%

Thanks